A series RC circuit has R=240 and C=102uF. It is connected in parallel to an inductor of 200mH and the combination is connected across a 220V, 60Hz source. Find the current in each branch.


1
Expert's answer
2021-11-24T19:17:02-0500

R=240 ohm

V=220V

"C=102\mu F"

f=60Hz

L=200mH

"Z=\sqrt{R^2+(X_L-X_c)^2}"


"X_L=wL=2\pi fL=2\pi\times60\times200\times10^{-3}=75.36\Omega"

"X_L=\frac{1}{2\pi fC}=\frac{1}{2\pi\times60\times102\times10^{-6}}=26.01\Omega"

"X_L-X_C=75.36-26.01=49.35\Omega"

Impedence

"Z=\sqrt{240^2+49.35^2}=245.02\Omega"

Current


"I=\frac{E}{Z}=\frac{220}{245.02}=0.8978Amp"

"Z_{RC}=\sqrt{R^2+X_C^2}"


"Z_{RC}=\sqrt{240^2+26.01^2}=241.405\Omega"

"Z_L=X_L"

Now current


"I_L=\frac{V}{X_L}"

"I_L=\frac{220}{75.36}=2.92Amp"

"I_{RC}=\frac{V}{Z_{RC}}=\frac{220}{241.36}=0.91Amp"